I have 2 different dish designs to buy. I need 250 sets. One design costs $20 and the other set $40. I have $6800 to spend. How many of each design can I order? x = number of $20 dishes and y = number of $40 dishes
What is the solution to this problem.

Answer by checkley77(12844)

X+Y=250 OR X=250-Y
20X+40Y=6,800
REPLACE X IN THE SECOND EQUATION WITH (25-Y) & SOLVE FOR Y.
20(250-Y+40Y=6,800
5,000-20Y+40Y=6,800
20Y=6,800-5,000
20Y=1,800
Y=1,800/20
Y=90 $40 DESIGN CAN BE BOUGHT.
250-90=160 $20 DESIGH CAN BE BOUGHT.
PROOF:
20*160+40*90=6,800
3,200+3,600=6,800
6,800=6,800